if i have the coordinates of a slope (-3,7)and (5,-1) how do i find the line of equation with the points
Found 2 solutions by stanbon, nerdybill:
Answer by stanbon(75887)

if i have the coordinates of a slope (-3,7)and (5,-1) how do i find the line of equation with the points
----------
slope = (-1-7) / (5--3) = -8/8 = -1
------------------
Equation is y = mx + b ; you have x,y and m ;solve for "b":
7 = -1*5 + b
b = 12
-------------
Equation:
y = -x + 12
================
Cheers,
Stan H.

Answer by nerdybill(7384)

First, find the slope from:
m = (y2-y1)/(x2-x1)
Given:(-3,7)and (5,-1)
.
m = (-1-7)/(5-(-3))
m = (-8)/(5+3)
m = (-8)/(8)
m = -1
.
Now, using the slope (-1) above and one of the two given points say (-3,7) plug it into the point slop form of:
(y-y1) = m(x-x1)
(y-7) = (-1)(x-(-3))
(y-7) = (-1)(x+3)
y-7 = -x-3
y = -x + 4 (this is what they're looking for in "slope intercept" form)